OpenStudy (anonymous):

Given the function f(x) = 2x +4 find the value of f(1/2). A. 5 B. 4 ½ C. 6 ½ D. 6

OpenStudy (anonymous):

Put \(x = \frac{1}{2}\) in the function..

jimthompson5910 (jim_thompson5910):

\[\Large f(x) = 2x + 4 \] \[\Large f\left( \frac{1}{2} \right ) = 2\left( \frac{1}{2} \right ) + 4 \]

jimthompson5910 (jim_thompson5910):

I'll let you finish

OpenStudy (anonymous):

After solving what you got @suckatmath17

OpenStudy (anonymous):

i really have no clue what to do...

OpenStudy (anonymous):

sorry

jimthompson5910 (jim_thompson5910):

2 times one-half is ???

OpenStudy (anonymous):

do i just times the bottom half of the fraction with 2?

jimthompson5910 (jim_thompson5910):

So \[\Large f(x) = 2x + 4 \] \[\Large f\left( \frac{1}{2} \right ) = 2\left( \frac{1}{2} \right ) + 4 \] \[\Large f\left( \frac{1}{2} \right ) = 1 + 4 \]

jimthompson5910 (jim_thompson5910):

sry made a typo

OpenStudy (anonymous):

im sorry i still dont get this at all!:(

jimthompson5910 (jim_thompson5910):

\[ \Large 2 \times \frac{1}{2} \] \[ \Large \frac{2}{1} \times \frac{1}{2} \] \[ \Large \frac{2 \times 1}{1 \times 2} \] \[ \Large \frac{2}{2} \] \[ \Large 1\] So \[ \Large 2 \times \frac{1}{2} = 1 \]

jimthompson5910 (jim_thompson5910):

Hopefully you now see how 2 times one-half is 1?

OpenStudy (anonymous):

yes i do thanks

jimthompson5910 (jim_thompson5910):

So this means... \[\Large f(x) = 2x + 4 \] \[\Large f\left( \frac{1}{2} \right ) = 2\left( \frac{1}{2} \right ) + 4 \] \[\Large f\left( \frac{1}{2} \right ) = 1 + 4 \] \[\Large f\left( \frac{1}{2} \right ) = 5 \]

OpenStudy (anonymous):

oh my gosh thanks soooo much

jimthompson5910 (jim_thompson5910):

you're welcome